OpenStudy (anonymous):

PLEASE HELP If the rate of inflation averages r per annum over n years, the amount A that $P will purchase after n years is A=P(1-r)^n, where r is expressed as a decimal. IF the average inflation rate is 2.2%, how long is it until the purchasing power is cut in half. It will be about____ years until the purchasing power is cut in half.

jimthompson5910 (jim_thompson5910):

A=P(1-r)^n P/2=P(1-r)^n 1/2=(1-r)^n 1/2=(1-0.022)^n 1/2 = 0.978^n keep going to solve for n

OpenStudy (anonymous):

do you do log to find n?

jimthompson5910 (jim_thompson5910):

take the log of both sides to get 1/2 = 0.978^n log( 1/2 ) = log( 0.978^n ) log( 1/2 ) = n*log( 0.978 ) log( 1/2 )/log( 0.978 ) = n n = log( 1/2 )/log( 0.978 ) n = ???

OpenStudy (anonymous):

thank you for taking you time to write all of this! very helpful!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

does 31.16 sound about right?

jimthompson5910 (jim_thompson5910):

31.1588314890099 which rounds to 31.16 so yes it does

jimthompson5910 (jim_thompson5910):

if you must do it to the nearest whole year, then it's 32 years (since you round up)

OpenStudy (anonymous):

thank you! can you help me with the one i posted up a while ago?

jimthompson5910 (jim_thompson5910):

post it here or send the link